sqrt(x)-sqrt(x-5)=1
Answer by mananth(16946)

sqrt(x)-sqrt(x-5)=1
x+(x-5)-2sqrtx*sqrt(x-5)=1
2x-5-1 = 2sqrt(x(x-5))
2x-6=2sqrt(x^2-5x)
/2
x-3=sqrt(x^2-5x)
square both sides
(x-3)^2=x^2-5x
x^2-6x+9=x^2-5x
6x-5x=9
x=9